OpenStudy (anonymous):

Solve: 3(x − 1) − (x + 5) = 2 3x-3-x+5=2 4x-3+5=2 4x-2=2 am I on the right track

OpenStudy (anonymous):

Not quite. This is where you made a mistake. \[3(x - 1) - (x + 5) = 2\]\[3x - 3 - x - 5 = 2\]Can you finish the rest?

OpenStudy (anonymous):

I got to the point.

OpenStudy (anonymous):

I actuall past that point

OpenStudy (anonymous):

Um. This is what you put: \[3x-3-x+5=2\]\[3x - 3 - x - 5 = 2\]Do you see the difference? The -1 is distributed.

OpenStudy (anonymous):

trying to make sure that I'm on the right path

OpenStudy (anonymous):

ok got you

OpenStudy (anonymous):

Alright. Tell me what you get for x afterwards so that I check it :)

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

sorry acutally X=1

OpenStudy (anonymous):

Not quite. \[3x - 3 - x - 5 = 2\]\[(3x - x) + (-3 - 5) = 2\]\[2x - 8 = 2\]\[2x = 10\]\[x = ?\]

OpenStudy (anonymous):

oh wow, not even close to what I was thing. I am always getting rid of parenthesis

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

np :)